OpenStudy (nuccioreggie):

The mass of a freight train is 9.98 ⋅ 106 kilograms. The mass of an aircraft carrier is 7.3 ⋅ 107 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

OpenStudy (nuccioreggie):

@mathmate

OpenStudy (nuccioreggie):

thank you for coming to helpme

OpenStudy (anonymous):

10^0=1 10^1=10 10^2=100 10^3=1,000 10^4=10,000 10^5=100,000 Therefore: 9 x 10^5 = 9 x 100,000 = 900,000 Answer: 9 x 10^5

OpenStudy (anonymous):

hoped i help

OpenStudy (nuccioreggie):

one more please and thank you

OpenStudy (anonymous):

sure

OpenStudy (nuccioreggie):

The sun produces 3.9 ⋅ 1033 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 3.25 ⋅ 103 seconds? (5 points) Part B: Which is the more reasonable measurement of the distance between the tracks on a railroad: 1.435 ⋅ 10−3 mm or 1.435 ⋅ 103 mm? Justify your answer. (5 points)

OpenStudy (nuccioreggie):

help whats the answer for part one and two

OpenStudy (anonymous):

@jayzdd